Smith, Vice Chancellor, Administration and Finance
Subject: Hurricane Katrina Recovery – Update #50
Posted: August 31, 2007
On August 23rd,
the State Fire Marshal granted a temporary (30 day)
occupancy permit for the Dental School Clinic and
Administration Buildings. There were a number of
small items identified during the inspection that
are being addressed so that the occupancy permit can
be made permanent. On Thursday, August 30, 2007, Mr.
George Thiaville of the State of Louisiana Office of
State Fire Marshall conducted a re-inspection of the
1100 Florida Avenue property as a follow-up to his
inspection the week before. The Fire Marshall (FM)
found that LSUHSC-NO had corrected all items noted
for correction on the August 23, 2007 inspection
report. Therefore, LSUHSC has been granted full
occupancy of the Clinic and Administration buildings
with the exception of the 1st floor.
The Dental School is
moving furniture, equipment, files, etc. back from
their temporary home on the LSU A&M South Campus in
Baton Rouge and they anticipate starting classes and
seeing patients in the near future. As was true for
the downtown campus buildings, the 1st
floors have not had their permanent repairs planned
or completed. The main entrance into the buildings
will be through the 1st floor of the
Clinic Building, where temporary repairs have been
made and a reception area is being prepared.
Work at the Dental School
also continues in a few areas. The Central
Sterilization unit is expected to be back in
operation in its new 3rd floor location
within the next week or so. Areas of the 2nd
floor of the Administration Building are being
carpeted. Auxiliary Enterprises will be establishing
a temporary food service area on the 4th
floor walkway between the buildings, and a temporary
Bookstore in a portion of the space currently
occupied by the Library. Temporary work spaces for
Facility Services, Supply Chain Management, and
Dental Central Supply are being constructed on the 1st
floor of the Clinic Building. The installation of
new ceilings on the 1st floor of the
Clinic is just getting underway. The restoration of
the exterior lighting in the parking lots and
grounds is substantially complete.
The potable water lines at
the Dental School Administration and Clinic
Buildings will be treated with chlorine starting
Friday, August 31st, from 8:30 am to Monday,
September 3rd at 1:00 pm. Contractors will be
shocking the water system and the water lines will
be filled with chlorine. As a result, the water will
be non-potable and students, faculty and staff are
urged not to drink the water or use it for washing
hands until Monday afternoon. The chlorine must sit
in the pipes and will be flushed on Monday morning,
Temporary repairs continue
to be made at 1542 Tulane Ave. The final stages of
work to restore power to the remaining portions of
the building are substantially complete. There
remain a few spot outages that are being researched
and repaired. The replacement of the flood damaged
and obsolete fire alarm system is progressing well.
The old devices have been removed and new wiring is
being pulled throughout the building as necessary.
The temporary emergency generator is on site and the
work on the construction of a foundation is getting
underway. A change proposal to repair the water
softener system has been approved and work will
begin on repairs shortly.
In our efforts to
continuously maintain and improve our facilities, we
are scheduling the painting of the Main Campus
elevated walkway. From 6:00 PM until 6:00 AM each
day for the period of Tuesday, September 4, 2007
through Wednesday, October 3, 2007, portions of the
walkway will be partially blocked off by our
contractor, Zimmer-Eschette. It is our plan to
maintain at least half of all passageways clear at
all times to ensure continuous access. The areas
affected will include the sections leading to the
Resource Center, South Roman Parking Garage, Allied
Health/School of Nursing, Lions/LSU Clinics,
Clinical Sciences, and Medical Education buildings.
The Walk to Wellness section will not be affected.
Here are highlights of
work that has been accomplished since our last
update was posted two weeks ago.
Project to Make Temporary
Repairs to the Dental School:
- The final environmental remediation and duct
cleaning contract is substantially complete.
- One air handler has required re-cleaning
- The project to upgrade the network
infrastructure is substantially complete.
- Doors for the new network/telecom
closets have not yet been delivered and
- The network upgrades have been completed
and the network is in operation.
- The telephone switch has been installed and
telephone instruments are being tested and
- The contract to restore the Central
Sterilization function in the 3rd floor location
in the Clinic Building space is nearing
- Both of the sterilizers were moved into
their final positions.
- The floor leveling is complete and the
new flooring has been installed.
- The installation of the last of the
peripheral equipment is substantially
- The system is expected to be operational
within the next week.
- The repairs to the exterior site lighting
and the electrical work necessary to power the
parking lot controls are substantially complete.
- The light poles and fixtures have been
delivered and most have been installed.
- The power has been restored to the
parking lot control devices.
- Carpet and floor tile is being installed on
the 2nd and 3rd floors of the Administration
- The carpeting of Auditorium C has been
- The installation of carpet in the Dean's
office is complete.
- The installation of new carpet on the
2nd floor of the Administration Building is
complete except for some threshold and
- A change order proposal to replace a number
of HVAC inline dampers in the buildings has been
received and approved.
- LSUHSC had a limited number of dampers
on site and several of the dampers have been
- Some of the old dampers have been found
to be sealed with tape that contains
asbestos. A small contract has been issued
for the removal of the asbestos tape so that
the remainder of the old dampers can be
- The rest of the dampers are on order and
are expected any day.
- The repairs to the compressed air line
running through the utility tunnel from the
Power Plant to the Clinic Building and the
related equipment has been completed.
- Representatives of the roof testing company
completed their analysis and provided the
results to the architects.
- The test results and the architects’
recommendation to replace the roofs on the
Clinic and Administration Building have been
forwarded to FEMA for their consideration.
- The architects are nearing completion of
the construction and bid documents for this
work. They are scheduled to be delivered
early next week.
- If FEMA concurs with the recommendation
to replace the roofs, bids for the project
can be advertised.
- The freight elevator is now operating in
manual mode with an operator.
- Work on the elevator disconnects has
- We are hopeful that the freight elevator
will be certified by the State Elevator
Inspector for normal use within the next
- Work on the Change Order to create temporary
work spaces for Supply Chain Management,
Facility Services and Central Supply on the 1st
floor of the Clinic Building has begun.
- Landis has leveled the floor in areas 1
- The final connections to the temporary
AHUs have been made.
- The walls in the elevator lobby are
being taped and floated and are ready for
- Electrical connections to outlets and
switches have been made in areas 1 and 2.
- The lights have been installed in area
- The fire panel annunciator was installed
near the temporary security desk on the 1st
- The walls in area 3 are being sheet
- Minor amounts of Asbestos identified in
the 1st floor mechanical room will be abated
this weekend allowing Landis to complete
connections for the air handling unit (AHU)
in area 3.
- The construction and installation of
temporary doors leading from the public
space on the 1st floor to areas 1
and 2 is complete.
- When the regular doors are
delivered, they will be installed and
the temporary doors removed.
- The construction of the fur down in the
elevator lobby is substantially complete.
- Electrical wiring continues to be run
throughout area 3.
- The construction of the mechanical room
in area 3 is nearly complete.
- Electrical equipment is being
- The damaged acid waste pipe located in the
ceiling area of the 1st floor of the Clinic
Building has been removed.
- The majority of piping has been
installed on the first floor with only a few
more connections needed.
- The new furniture for the Clinic waiting
areas arrived and is being assembled.
Project to Make Temporary
Repairs to 1542 Tulane Ave.:
- Ordes successfully completed the scheduled
shut-down and installation of the current
transformers and amp meter of sub-station PL-1.
- Three of the remaining 4th floor riser
fuses have been delivered to the site and
- Ordes is making preparations for the
installation of the emergency generator.
- Ordes removed their storage container
from the site to make room for the
construction and installation of the
concrete foundation for the temporary
- The layout for the foundation is
- Siemens is still removing the obsolete fire
alarm devices and the related wiring on the
upper floors of the building.
- They have installed a new riser for the
fire alarm wiring and is pulling wire
through the new riser.
- They have been adding junction boxes on
each floor in the new fire alarm riser.
- Siemens is routing conduit from the new
riser through the basement to the fire alarm
cabinets on the 1st floor.
- Stratos continues their repair work on
- They installed the tape selector and the
TLS box which controls the elevator leveling
for Elevator #2.
- They are terminating the wiring of the
travel cables for Elevator #2.
- Stratos is replacing damaged equipment
on top of the elevator car.
- Stratos re-programmed the phone line for
- Stratos is also working on the installation
of the controllers for Elevators #7 and #8,
installing new conduit and wiring connections.
- The new cabs for Elevators #7 & #8 have
- Ordes delivered and installed the interim
emergency generator secured by LSUHSC to serve
until the temporary emergency generator
foundation is build and the generator is mounted
- This interim generator will
automatically power the basement sump pumps
in the event of a power outage.
- It is connected to the ATS and will
energize when the city power is lost.
- The contractor also connected power to gate
controllers at the L & M Lot and the Lakeside
- The crews cleaned and graded the new
planting beds on the S. Prieur St. side of the
Lions/LSU Clinic Building.
- Four new Savannah holly trees were planted
on the S. Prieur St. side of the Lions/LSU
- The crews continued planting on the S.
Prieur St. side of the Lions/LSU Clinic Building
and completed the planting of the remaining
bushes. This completes the planting around the
- Crews sprayed turf areas on the Gravier St.
side of the Allied Health/Nursing Building with
herbicide to kill the existing vegetation as
they prepare to create new planting beds.
- Crews also worked on preparing trenches on
Bolivar St. side of the MEB and the Allied
Health/Nursing Building to run the wire for
controlling the irrigation system.
MEB Cafeteria Project:
- Most of the hanging light fixtures have been
- Lights around the perimeter where
booths, etc. will be installed beneath the
fixtures locations will not be installed
until the furniture has been delivered and
set in place.
- The frames for the curved glass wall along
the hallway have been installed.
- All of the glass panels have been
installed in the frames.
- A new Starbucks Coffee kiosk is being
assembled in the atrium between the Allied
Health/Nursing Building and the MEB.
MEB Duct Cleaning:
- The cleaning of the 5th floor ducts is
- The cleaning of the 6th floor
ducts is complete.
- The cleaning of the 7th floor
ducts is complete.
- The project is complete.
MEB Morgue Upgrade:
- The original morgue cooler in the MEB had
not been upgraded since the building was
constructed. Significant changes and
enhancements within the industry have come about
since that time, so a new cooler and related
storage has been purchased.
- The old cooler has been demolished to
make way for the new cooler installation.
- The new concrete floor has been poured.
- The new insulated panels that will make
up the walls and ceiling of the cooler are
scheduled to be delivered next week,
- The current target date for completion
of the work is in early October.
- A project is underway to replace all of the
old lights in the Roman St. Garage with new
lights that are both brighter and more energy
efficient. The result will be brighter light
throughout the parking garage, without any
increase in energy consumption.
- Lights have been replaced throughout the
entire garage structure.
We encourage you to follow the
details of our progress through the daily updates
which are posted on the web site.